Circular No. 6194

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) The following improved parabolic orbital elements, from MPEC 1995-

P01, are based on 208 observations July 24-Aug. 1:

T = 1997 Apr. 1.810 TT  ω = 129.^{o}956
Ω = 282.339 }2000.0
q = 0.92819 AU  i = 89.779

These elements are still somewhat uncertain but indicate minimum approach distances of 0.75 AU to Jupiter on 1996 Apr. 5 and 1.32 AU to the earth on 1997 Mar. 23.

Total visual magnitude estimates: July 26.22 UT, 10.8 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 27.19, 10.7 (Hale); 28.15, 10.7 (Hale); 28.24, 9.9 (R. Keen, Mt. Thorodin, CO, 0.08-m reflector); 28.94, 10.3 (J. Carvajal, Madrid, Spain, 0.44-m reflector); 29.31, 10.5 (C. S. Morris, Lockwood Valley, CA, 0.26-m reflector); 29.93, 10.2, (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 30.28, 10.5 (C. S. Morris, Pine Mountain Club, CA, 0.26-m reflector); 30.29, 10.9 (Hale); 30.96, 10.4 (Carvajal). Prediscovery observations—especially negative prediscovery observations—of the comet would be useful.

W. Offutt, Cloudcroft, NM, reports that his CCD imaging with a 0.6m reflector suggests that the northward “tail” mentioned on IAUC 6188 appears to be the maximum extension of an asymmetric coma and that, although not all of the images were identically exposed, the coma seems to have been shrinking over the course of the past week. The position angle of this maximum extension, its extent from the nucleus, its extent on the opposite side of the nucleus, and the coma width perpendicular to the extension, are as follows: July 24.19 UT, 14^{o}, 154′′, 81′′, 218′′; 24.33, 9^{o}, 154′′, 51′′, 202′′; 25.15, 23^{o}, 149′′, 55′′, 179′′; 25.26, 13^{o}, 138′′, 53′′, 188′′; 28.18, 13^{o}, 138′′, 41′′, 180′′; 29.14, 2^{o}, 109′′, 38′′, 172′′; 31.19, 10^{o}, 111′′, 39′′, 161′′. In all cases, the edge of the coma is least clearly defined near the position angle of maximum extension and is most clearly defined in the opposite direction. Commenting on the above, Z. Sekanina, Jet Propulsion Laboratory, suggests that the observations are of a spiral coma, similar to the halo sometimes displayed by comet 29P/Schwassmann-Wachmann 1 after an outburst, the shrinking and the diffuseness in the direction of maximum extension apparently products of a rapid fading of the halo’s outer limits. There seems little question that comet C/1995 O1 has had an outburst, and it may grow fainter for some time.

1995 August 1 Brian G. Marsden